# On the Identifying Power of Generalized Monotonicity

### v3

@@ Abstract
In the context of a binary outcome, treatment, and instrument, Balke and Pearl (1993, 1997) establish that the monotonicity condition of Imbens and Angrist (1994) has no identifying power beyond instrument exogeneity for average potential outcomes and average treatment effects in the sense that adding it to instrument exogeneity does not decrease the identified sets for those parameters whenever those restrictions are consistent with the distribution of the observable data. This paper shows that this phenomenon holds in a broader setting with a multi-valued outcome, treatment, and instrument, under an extension of the monotonicity condition that we refer to as generalized monotonicity. We further show that this phenomenon holds for any restriction on treatment response that is stronger than generalized monotonicity provided that these stronger restrictions do not restrict potential outcomes. Importantly, many models of potential treatments previously considered in the literature imply generalized monotonicity, including the types of monotonicity restrictions considered by Kline and Walters (2016), Kirkeboen et al. (2016), and Heckman and Pinto (2018), and the restriction that treatment selection is determined by particular classes of additive random utility models. We show through a series of examples that restrictions on potential treatments can provide identifying power beyond instrument exogeneity for average potential outcomes and average treatment effects when the restrictions imply that the generalized monotonicity condition is violated. In this way, our results shed light on the types of restrictions required for help in identifying average potential outcomes and average treatment effects.
@@ 1 Introduction
In their analysis of a setting with a binary outcome, treatment, and instrument, Balke and Pearl (1993, 1997) establish that the monotonicity condition of Imbens and Angrist (1994) has no identifying power beyond instrument exogeneity for average potential outcomes and the average treatment effect (ATE). Here, by no identifying power beyond instrument exogeneity, we mean that adding the monotonicity condition of Imbens and Angrist (1994) to instrument exogeneity does not decrease the identified sets for those parameters whenever those restrictions are consistent with the distribution of the observable data. [^1] For settings with possibly non-binary outcomes, Kitagawa (2021) shows this phenomenon continues to hold for any parameter that is a function of the marginal distributions of potential outcomes. In this way, their results contrast with the analysis of Imbens and Angrist (1994), who showed that their monotonicity condition and instrument exogeneity permitted identification of the local average treatment effect (LATE). This paper studies the extent to which this phenomenon holds in the broader context of a multi-valued outcome, treatment, and instrument.
@@
We show that a generalization of the monotonicity condition of Imbens and Angrist (1994) to this richer setting also has no identifying power beyond instrument exogeneity for average potential outcomes and ATEs. We hereafter refer to this condition more succinctly as generalized monotonicity. We show further that this result remains true for any restriction that is in fact stronger than generalized monotonicity provided that these stronger restrictions do not restrict potential outcomes in a sense that we will make precise later. This feature of our results is remarkable because one might expect stronger restrictions, possibly with very complicated restrictions on potential treatments that are difficult to fully characterize, to reduce the identified set at least in some instances, but we show that this is not the case. Using this remarkable result, our analysis accommodates many examples of restrictions on potential treatments that have been previously considered in the literature. In particular, we show that encouragement designs, the types of monotonicity restrictions considered by Kline and Walters (2016), Kirkeboen et al. (2016), and Heckman and Pinto (2018), and certain additive random utility models, including some studied in Lee and Salanié (2023), all satisfy generalized monotonicity. We further establish through a series of examples that models that do not satisfy generalized monotonicity can in fact have identifying power for average potential outcomes and ATEs. In this way, our analysis sheds light on the types of restrictions required for help in identifying average potential outcomes and ATEs.
@@
In establishing our results, we derive the identified sets for average potential outcomes under any such restriction and instrument exogeneity while maintaining the assumption that these restrictions are consistent with the distribution of the observable data. Our derivations reveal that the form of the resulting identified sets parallels the form of those derived by Balke and Pearl (1993, 1997) for a binary outcome, treatment, and instrument. An implication of the form of the identified sets is that average potential outcomes and ATEs are only identified under an identification-at-infinity-type condition when imposing instrument exogeneity and any such restriction. In our analysis, we also derive the identified sets for average potential outcomes and ATEs when imposing instrument exogeneity alone whenever the distribution of the observable data is consistent with instrument exogeneity and generalized monotonicity. As we explain in Example 4.1, this consistency is necessarily satisfied, for example, in the context of a multi-arm randomized controlled trial with one-sided non-compliance when defining the instrument to be random assignment to a given treatment arm.
@@
Our paper differs from the closely related literature that, in the context of a binary outcome, treatment, and instrument, considers the identifying power of the monotonicity condition of Imbens and Angrist (1994) and instrument exogeneity for the distribution (as opposed to the average) of potential outcomes, or considers the identifying power of these conditions when combined with additional restrictions on potential outcomes. In particular, Kamat (2019) shows that the monotonicity condition of Imbens and Angrist (1994) does have identifying power beyond instrument exogeneity for the (joint) distribution of potential outcomes. Machado et al. (2019) show that the monotonicity condition of Imbens and Angrist (1994) does have additional identifying power for the ATE beyond instrument exogeneity if one additionally imposes an assumption that requires potential outcomes to vary monotonically with the treatment. Thus, the phenomenon we explore is sensitive to both the choice of parameter and to whether one imposes assumptions on potential outcomes.
@@ 2 Setup and Notation
i.e., the the pre-image of $P$ under $T$. We say $\mathbf{Q}$ is consistent with $P$ if and only if $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. For every model $\mathbf{Q}$ considered in this paper, every $Q\in\mathbf{Q}$ is assumed to satisfy:
@@
See Richardson and Robins (2013) for an analysis of alternative exogeneity assumptions, and in particular how this exogeneity assumption avoids the “cross-world” restrictions of the stronger joint independence assumption $((Y_{d}:d\in\mathcal{D}),(D_{z}:z\in\mathcal{Z}))\perp\!\!\!\perp Z$ under $Q$. Under this assumption, $\mathbf{Q}_{0}(P,\mathbf{Q})$ can be simplified as follows. If $Q$ satisfies Assumption 2.1, then
@@
$P=QT^{-1}$ if and only if (2) holds. Thus, if all $Q\in\mathbf{Q}$ satisfies Assumption 2.1, then
$\mathbf{Q}_{0}(P,\mathbf{Q})=\bigg{\{}Q\in\mathbf{Q}:~{}p_{yd|z}=Q\{Y_{d}=y,D_{z}=d\}\text{ for all }y\in\mathcal{Y},d\in\mathcal{D},z\in\mathcal{Z}\bigg{\}}~{}.$
@@
In what follows, we refer to Assumption 2.2 as generalized monotonicity. It states that under $Q$, for each treatment status $d\in\mathcal{D}$, there exists a value (possibly depending on $d$ and $Q$) of the instrument $z^{\ast}\in\mathcal{Z}$ that maximally encourages all individuals to $d$. Here, by “maximally encourage”, we mean that if an individual does not choose $d$ when $Z=z^{\ast}$, then they never choose $d$ for any other value of $Z$. Equivalently, if an individual chooses $d$ when $Z$ is equal to any value other than $z^{\ast}$, then they have to choose $d$ when $Z=z^{\ast}$. When $\mathcal{D}=\mathcal{Z}=\{0,1\}$, Assumption 2.2 is equivalent to the monotonicity assumption of Imbens and Angrist (1994).
We emphasize that Assumption 2.2 only requires, for each possible value of the treatment, that there exists a value of the instrument that maximally encourages that treatment; it does not require that the value of the instrument is unique. For a given distribution $Q$ and given treatment $d\in\mathcal{D}$, let $\mathcal{Z}^{\ast}(d,Q)$ denote the set of $z^{\ast}$ that satisfy (3). In this notation, Assumption 2.2 can be restated as $\mathcal{Z}^{\ast}(d,Q)\neq\emptyset$ for each $d\in\mathcal{D}$. In the statement of Assumption 2.2, $z^{\ast}(d,Q)$ is allowed to change across $Q$. The following lemma shows that $\mathcal{Z}^{\ast}(d,Q)$ is identified from $P$ and is hence the same for all $Q$ that rationalizes $P$ and satisfies Assumptions 2.1 and 2.2. In what follows, we will therefore write $\mathcal{Z}^{\ast}(d)$ and $z^{\ast}(d)$ whenever the given distribution $Q$ rationalizes $P$ and satisfies Assumptions 2.1 and 2.2. This result generalizes the corresponding result in Imbens and Angrist (1994).
@@ 3 Main Result
In order to describe our main result, we first introduce some further notation. Denote by $\mathbf{Q}_{E}^{\ast}$ (where $E$ stands for exogeneity) the set of all distributions that satisfy Assumption 2.1 and by $\mathbf{Q}_{E,M}^{\ast}$ (where $M$ stands for generalized monotonicity) the set of all distributions that satisfy Assumptions 2.1 and 2.2. Denote by $Q|_{Y}$ the marginal distribution of $(Y_{d}:d\in\mathcal{D})$ under $Q$ and define $\mathbf{Q}|_{Y}=\{Q|_{Y}:Q\in\mathbf{Q}\}$. We will further require that the model does not restrict potential outcomes in the following sense:
@@
$\mathbf{Q}|_{Y}$ equals the entire set of distributions on $\mathcal{Y}^{|\mathcal{D}|}$.
@@
Suppose $\mathbf{Q}\subseteq\mathbf{Q}_{E,M}^{\ast}$ and $\mathbf{Q}|_{Y}$ satisfies Assumption 3.1. Then, for any $P$ such that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$, we have $\Theta_{0}(P,\mathbf{Q})=\Theta_{0}(P,\mathbf{Q}_{E,M}^{\ast})=\Theta_{0}(P,\mathbf{Q}_{E}^{\ast})$.
Theorem 3.1 describes the sense in which restrictions on potential treatments stronger than generalized monotonicity have no identifying power for average potential outcomes and ATEs provided that these stronger restrictions do not restrict potential outcomes. This result is established through Theorems 3.2 and 3.3 below. Theorem 3.2, developed in Section 3.1, characterizes $\Theta_{0}(P,\mathbf{Q})$ for any model $\mathbf{Q}$ that is stronger than Assumptions 2.1 and 2.2, i.e., $\mathbf{Q}\subseteq\mathbf{Q}_{E,M}^{\ast}$, and does not restrict potential outcomes in the sense of Assumption 3.1. The result shows, in particular, that $\Theta_{0}(P,\mathbf{Q})=\Theta_{0}(P,\mathbf{Q}_{E,M}^{\ast})$ for any such model $\mathbf{Q}$ whenever $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. Remarkably, this result holds even if the model $\mathbf{Q}$ is strictly more restrictive than Assumptions 2.1 and 2.2 in the sense that $\mathbf{Q}\subsetneqq\mathbf{Q}_{E,M}^{\ast}$. On the other hand, Theorem 3.3, developed in Section 3.2, shows that $\Theta_{0}(P,\mathbf{Q}_{E,M}^{\ast})=\Theta_{0}(P,\mathbf{Q}_{E}^{\ast})$ whenever $\mathbf{Q}_{0}(P,\mathbf{Q}_{E,M}^{\ast})\neq\emptyset$. Together, these results immediately imply Theorem 3.1.
@@ 3.1 Identified Sets for $\mathbf{Q}\subseteq\mathbf{Q}_{E,M}^{\ast}$
Suppose $\mathbf{Q}\subseteq\mathbf{Q}_{E,M}^{\ast}$ and $\mathbf{Q}|_{Y}$ satisfies Assumption 3.1. Then, for any $P$ such that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$,
@@
We now describe some intuition for Theorem 3.2. Note that the distribution of the data, $P$, only contains information on the distribution of $Y_{d}$ for those individuals who would take treatment $d$ for some value of the instrument. On the other hand, it contains no information on the distribution of $Y_{d}$ for those individuals who would not take that treatment for any value of the instrument. Assumption 2.2 implies that individuals would take treatment $d$ at some value of the instrument if and only if $D_{z^{\ast}(d)}=d$, i.e., when maximally encouraged to do so. Assumption 2.1 implies $\beta_{d|z^{\ast}(d)}=\mathbb{E}_{Q}[\mathbbm{1}\{D_{z^{\ast}(d)}=d\}Y_{d}]$, and thus captures all the information from $P$ relevant to $\mathbb{E}_{Q}[Y_{d}]$, which is the first part of the lower and upper bounds in (5). In contrast, Assumptions 2.1 and 2.2 imply that the probability that an individual would not take treatment $d$ for any value of $Z$ is identified from $P$ to be $Q\{D_{z^{\ast}(d)}\neq d\}=1-\sum_{y\in\mathcal{Y}}p_{yd|z^{\ast}(d)},$ but $P$ contains no information on the distribution of $Y_{d}$ for such individuals. Furthermore, that $\mathbf{Q}|_{Y}$ satisfies Assumption 3.1 implies that the model does not restrict the distribution of $Y_{d}$ for such individuals beyond $y\in\mathcal{Y}$, so that we can set $Y_{d}$ to be any value between $y^{L}$ and $y^{U}$ for these individuals, which constitutes the second part of the upper and lower bounds in (5).
@@
Under the instrument exogeneity and monotonicity assumptions of Imbens and Angrist (1994), Balke and Pearl (1993, 1997) found the same form of the identified set for $\theta(Q)$ as (5) when $\mathcal{Y}=\mathcal{D}=\mathcal{Z}=\{0,1\}$. Theorem 3.2 therefore generalizes the result of Balke and Pearl (1993, 1997) to more than two treatment arms and instrument values, to outcomes taking more than two values, and, more surprisingly, to show that the same identified set holds when imposing possibly stronger restrictions on potential treatments than generalized monotonicity.
@@
Therefore, when imposing generalized monotonicity, as well as when imposing any restriction implying generalized monotonicity, the ATE of $j$ versus $k$ is only identified “at infinity” (Heckman, 1990; Andrews and Schafgans, 1998) in the sense that identification requires

### v4

@@ Abstract
In the context of a binary outcome, treatment, and instrument, balke1993nonparametric; balke1997bounds establish that the monotonicity condition of imbens1994identification has no identifying power beyond instrument exogeneity for average potential outcomes and average treatment effects in the sense that adding it to instrument exogeneity does not decrease the identified sets for those parameters whenever those restrictions are consistent with the distribution of the observable data. This paper shows that this phenomenon holds in a broader setting with a multi-valued outcome, treatment, and instrument, under an extension of the monotonicity condition that we refer to as generalized monotonicity. We further show that this phenomenon holds for any restriction on treatment response that is stronger than generalized monotonicity provided that these stronger restrictions do not restrict potential outcomes. Importantly, many models of potential treatments previously considered in the literature imply generalized monotonicity, including the types of monotonicity restrictions considered by kline2016evaluating, kirkeboen2016field, and heckman2018unordered, and the restriction that treatment selection is determined by particular classes of additive random utility models. We show through a series of examples that restrictions on potential treatments can provide identifying power beyond instrument exogeneity for average potential outcomes and average treatment effects when the restrictions imply that the generalized monotonicity condition is violated. In this way, our results shed light on the types of restrictions required for help in identifying average potential outcomes and average treatment effects.
@@ 1 Introduction
In their analysis of a setting with a binary outcome, treatment, and instrument, balke1993nonparametric; balke1997bounds establish that the monotonicity condition of imbens1994identification has no identifying power beyond instrument exogeneity for average potential outcomes and the average treatment effect (ATE). Here, by no identifying power beyond instrument exogeneity, we mean that adding the monotonicity condition of imbens1994identification to instrument exogeneity does not decrease the identified sets for those parameters whenever those restrictions are consistent with the distribution of the observable data. [^1] For settings with possibly non-binary outcomes, kitagawa2021identification shows this phenomenon continues to hold for any parameter that is a function of the marginal distributions of potential outcomes. In this way, their results contrast with the analysis of imbens1994identification, who showed that their monotonicity condition and instrument exogeneity permitted identification of the local average treatment effect (LATE). This paper studies the extent to which this phenomenon holds in the broader context of a multi-valued outcome, treatment, and instrument.
@@
We show that a generalization of the monotonicity condition of imbens1994identification to this richer setting also has no identifying power beyond instrument exogeneity for average potential outcomes and ATEs. We hereafter refer to this condition more succinctly as generalized monotonicity. We show further that this result remains true for any restriction that is in fact stronger than generalized monotonicity provided that these stronger restrictions do not restrict potential outcomes in a sense that we will make precise later. This feature of our results is remarkable because one might expect stronger restrictions, possibly with very complicated restrictions on potential treatments that are difficult to fully characterize, to reduce the identified set at least in some instances, but we show that this is not the case. Using this result, our analysis accommodates many examples of restrictions on potential treatments that have been previously considered in the literature. In particular, we show that encouragement designs, the types of monotonicity restrictions considered by kline2016evaluating, kirkeboen2016field, and heckman2018unordered, and certain additive random utility models, including some studied in lee2023treatment, all satisfy generalized monotonicity.
@@
In establishing our results, we derive the identified sets for average potential outcomes under any such restriction and instrument exogeneity while maintaining the assumption that these restrictions are consistent with the distribution of the observable data. Our derivations reveal that the form of the resulting identified sets parallels the form of those derived by balke1993nonparametric; balke1997bounds for a binary outcome, treatment, and instrument. An implication of the form of the identified sets is that average potential outcomes and ATEs are only identified under an identification-at-infinity-type condition when imposing instrument exogeneity and any such restriction. In our analysis, we also derive the identified sets for average potential outcomes and ATEs when imposing instrument exogeneity alone whenever the distribution of the observable data is consistent with instrument exogeneity and generalized monotonicity. As we explain in Example 4.1, this consistency is necessarily satisfied, for example, in the context of a multi-arm randomized controlled trial with one-sided non-compliance when defining the instrument to be random assignment to a given treatment arm.
Our results further provide necessary conditions on restrictions on potential treatments to help in identifying average potential outcomes and ATEs. See Theorem 3.3 and the subsequent discussion for details. We illustrate this phenomenon through a series of examples of models that need not satisfy generalized monotonicity and have identifying power for average potential outcomes and ATEs.
@@
Our paper differs from the closely related literature that, in the context of a binary outcome, treatment, and instrument, considers the identifying power of the monotonicity condition of imbens1994identification and instrument exogeneity for the distribution (as opposed to the average) of potential outcomes, or considers the identifying power of these conditions when combined with additional restrictions on potential outcomes. In particular, kamat2019identifying shows that the monotonicity condition of imbens1994identification does have identifying power beyond instrument exogeneity for the (joint) distribution of potential outcomes. machado2019instrumental show that the monotonicity condition of imbens1994identification does have additional identifying power for the ATE beyond instrument exogeneity if one additionally imposes an assumption that requires potential outcomes to vary monotonically with the treatment. Thus, the phenomenon we explore is sensitive to both the choice of parameter and to whether one imposes assumptions on potential outcomes.
@@ 2 Setup and Notation
i.e., the pre-image of $P$ under $T$. We say $\mathbf{Q}$ is consistent with $P$ if and only if $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. We will start by considering models $\mathbf{Q}$ for which every $Q\in\mathbf{Q}$ satisfies
@@
$((Y_{d}:d\in\mathcal{D}),(D_{z}:z\in\mathcal{Z}))\perp\!\!\!\perp Z$ under $Q$.
Our final result on the identifying power of generalized monotonicity will also apply to the weaker exogeneity restriction in richardson2013single that avoids the “cross-world” restrictions of Assumption 2.1; see Assumption 3.3 and Corollary 3.3 below in Section 3.2. If $Q$ satisfies Assumption 2.1, then
@@
$P=QT^{-1}$ if and only if (2) holds. Thus, if all $Q\in\mathbf{Q}$ satisfies Assumption 2.1, then $\mathbf{Q}_{0}(P,\mathbf{Q})$ can be simplified as
$\mathbf{Q}_{0}(P,\mathbf{Q})=\bigg\{Q\in\mathbf{Q}:~p_{yd|z}=Q\{Y_{d}=y,D_{z}=d\}\text{ for all }y\in\mathcal{Y},d\in\mathcal{D},z\in\mathcal{Z}\bigg\}~.$ (3)
@@
In what follows, we refer to Assumption 2.2 as generalized monotonicity. It states that under $Q$, for each treatment status $d\in\mathcal{D}$, there exists a value (possibly depending on $d$ and $Q$) of the instrument $z^{\ast}\in\mathcal{Z}$ that maximally encourages all individuals to $d$. Here, by “maximally encourage”, we mean that if an individual does not choose $d$ when $Z=z^{\ast}$, then they never choose $d$ for any other value of $Z$. Equivalently, if an individual chooses $d$ when $Z$ is equal to any value other than $z^{\ast}$, then they have to choose $d$ when $Z=z^{\ast}$. When $\mathcal{D}=\mathcal{Z}=\{0,1\}$, Assumption 2.2 is equivalent to the monotonicity assumption of imbens1994identification.
We emphasize that Assumption 2.2 only requires, for each possible value of the treatment, that there exists a value of the instrument that maximally encourages that treatment; it does not require that the value of the instrument is unique. For a given distribution $Q$ and given treatment $d\in\mathcal{D}$, let $\mathcal{Z}^{\ast}(d,Q)$ denote the set of $z^{\ast}$ that satisfy (4). In this notation, Assumption 2.2 can be restated as $\mathcal{Z}^{\ast}(d,Q)\neq\emptyset$ for each $d\in\mathcal{D}$. In the statement of Assumption 2.2, $z^{\ast}(d,Q)$ is allowed to change across $Q$. The following lemma shows that $\mathcal{Z}^{\ast}(d,Q)$ is identified from $P$ and is hence the same for all $Q$ that rationalizes $P$ and satisfies Assumptions 2.1 and 2.2. In what follows, we will therefore write $\mathcal{Z}^{\ast}(d)$ and $z^{\ast}(d)$ whenever the given distribution $Q$ rationalizes $P$ and satisfies Assumptions 2.1 and 2.2. This result generalizes the corresponding result in imbens1994identification.
@@ 3 Main Result
In order to describe our main result, we first introduce some further notation. Denote by $\mathbf{Q}_{E}^{\ast}$ (where $E$ stands for exogeneity) the set of all distributions that satisfy Assumption 2.1 and by $\mathbf{Q}_{E,M}^{\ast}$ (where $M$ stands for generalized monotonicity) the set of all distributions that satisfy Assumptions 2.1 and 2.2. We will further require that the model does not restrict potential outcomes in the following sense:
@@
Let $Q\in\mathbf{Q}$ and $Q^{\prime}\in\mathbf{Q}_{E}^{\ast}$. If the distributions of $(D_{z}:z\in\mathcal{Z})$ under $Q$ and $Q^{\prime}$ are the same, then $Q^{\prime}\in\mathbf{Q}$.
@@
Suppose $\mathbf{Q}\subseteq\mathbf{Q}_{E,M}^{\ast}$ and $\mathbf{Q}$ satisfies Assumption 3.1. Then, for any $P$ such that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$, we have $\Theta_{0}(P,\mathbf{Q})=\Theta_{0}(P,\mathbf{Q}_{E,M}^{\ast})=\Theta_{0}(P,\mathbf{Q}_{E}^{\ast})$.
Theorem 3.1 describes the sense in which restrictions on potential treatments stronger than generalized monotonicity have no identifying power for average potential outcomes and ATEs provided that these stronger restrictions do not restrict potential outcomes. This result is established through Theorems 3.2 and 3.3 below. Theorem 3.2, developed in Section 3.1, characterizes $\Theta_{0}(P,\mathbf{Q})$ for any model $\mathbf{Q}$ that is stronger than Assumptions 2.1 and 2.2, i.e., $\mathbf{Q}\subseteq\mathbf{Q}_{E,M}^{\ast}$, and does not restrict potential outcomes in the sense of Assumption 3.1. The result shows, in particular, that $\Theta_{0}(P,\mathbf{Q})=\Theta_{0}(P,\mathbf{Q}_{E,M}^{\ast})$ for any such model $\mathbf{Q}$ whenever $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. Remarkably, this result holds even if the model $\mathbf{Q}$ is strictly more restrictive than Assumptions 2.1 and 2.2 in the sense that $\mathbf{Q}\subsetneqq\mathbf{Q}_{E,M}^{\ast}$. On the other hand, Theorem 3.3 and Corollary 3.2, developed in Section 3.2, show that $\Theta_{0}(P,\mathbf{Q}_{E,M}^{\ast})=\Theta_{0}(P,\mathbf{Q}_{E}^{\ast})$ whenever $\mathbf{Q}_{0}(P,\mathbf{Q}_{E,M}^{\ast})\neq\emptyset$. Together, these results immediately imply Theorem 3.1. In fact, Theorem 3.3 shows the stronger result that if a submodel of instrument exogeneity and generalized monotonicity is consistent with $P$, then any model sandwiched between this submodel and the model that only assumes mean independence leads to the same identified set for average potential outcomes and ATEs. This observation allows us to establish that generalized monotonicity also has no identifying power for average potential outcomes and ATEs beyond the weaker exogeneity restriction of richardson2013single.
@@ 3.1 Identified Sets for $\mathbf{Q}\subseteq\mathbf{Q}_{E,M}^{\ast}$
Suppose $\mathbf{Q}\subseteq\mathbf{Q}_{E,M}^{\ast}$ and $\mathbf{Q}$ satisfies Assumption 3.1. Then, for any $P$ such that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$,
@@
We now describe some intuition for Theorem 3.2. Note that the distribution of the data, $P$, only contains information on the distribution of $Y_{d}$ for those individuals who would take treatment $d$ for some value of the instrument. On the other hand, it contains no information on the distribution of $Y_{d}$ for those individuals who would not take that treatment for any value of the instrument. Assumption 2.2 implies that individuals would take treatment $d$ at some value of the instrument if and only if $D_{z^{\ast}(d)}=d$, i.e., when maximally encouraged to do so. Assumption 2.1 implies $\beta_{d|z^{\ast}(d)}=\mathbb{E}_{Q}[\mathbbm{1}\{D_{z^{\ast}(d)}=d\}Y_{d}]$, and thus captures all the information from $P$ relevant to $\mathbb{E}_{Q}[Y_{d}]$, which is the first part of the lower and upper bounds in (6). In contrast, Assumptions 2.1 and 2.2 imply that the probability that an individual would not take treatment $d$ for any value of $Z$ is identified from $P$ to be $Q\{D_{z^{\ast}(d)}\neq d\}=1-\sum_{y\in\mathcal{Y}}p_{yd|z^{\ast}(d)},$ but $P$ contains no information on the distribution of $Y_{d}$ for such individuals. Furthermore, that $\mathbf{Q}$ satisfies Assumption 3.1 implies that the model does not restrict the distribution of $Y_{d}$ for such individuals beyond $y\in\mathcal{Y}$, so that we can set $Y_{d}$ to be any value between $y^{L}$ and $y^{U}$ for these individuals, which constitutes the second part of the upper and lower bounds in (6).
@@
Under the instrument exogeneity and monotonicity assumptions of imbens1994identification, balke1993nonparametric; balke1997bounds found the same form of the identified set for $\theta(Q)$ as (6) when $\mathcal{Y}=\mathcal{D}=\mathcal{Z}=\{0,1\}$. Theorem 3.2 therefore generalizes the result of balke1993nonparametric; balke1997bounds to more than two treatment arms and instrument values, to outcomes taking more than two values, and, more surprisingly, to show that the same identified set holds when imposing possibly stronger restrictions on potential treatments than generalized monotonicity.
@@
Therefore, when imposing generalized monotonicity, as well as when imposing any restriction implying generalized monotonicity, the ATE of $j$ versus $k$ is only identified “at infinity” (heckman1990varieties; andrews1998semiparametric) in the sense that identification requires